# Structure constants of Peterson Schubert calculus

### v1

@@ 2.3. Cohomology ring of Peterson varieties and Peterson Schubert classes
Consider the restriction of the line bundle $L_{\mu}$ on the Peterson variety $\operatorname{Pet}_{G}$, which will also be denoted by $L_{\mu}$ when there is no confusion. Harada–Horiguchi–Masuda [19] gave the following explicit and beautiful Borel-type presentation of the equivariant cohomology ring of Peterson varieties in all Lie types.
@@
Suppose that rank of $G$ is $n$. The cohomology ring of the Peterson variety $\operatorname{Pet}_{G}$ has the following presentation
@@
It follows that while class $\iota^{*}\left(\sigma_{v_{I}}\right)$ depends on the choice of $v_{I}$, the class $p_{I}$ is independent of this choice. In [16, Remark 7.7], a type-independent formula for the intersection multiplicity $m\left(v_{I}\right)$ was conjectured, which was proved in [17].
@@
Next, we recall the equivariant Giambelli formula for the Peterson variety in [13] and [17]. It is well known that equivariant Schubert classes $\sigma_{s_{i}},i\in\Delta$, generate $H_{T}^{*}(G/B;\mathbb{Q})$ as a $H_{T}^{*}(pt)$-algebra. Since the restriction map $H_{T}^{*}\left(G/B;\mathbb{Q}\right)\rightarrow H_{S}^{*}\left(\operatorname{Pet}_{G};\mathbb{Q}\right)$ is surjective [19], it is easy to see that each $\iota^{*}\sigma_{v_{I}}$ can be expressed as a polynomial with $\mathbb{Q}[t]$-coefficients in $\iota^{*}\sigma_{s_{i}},i\in\Delta$. Drellich [13] gave a very simple equivariant Giambelli formula for this polynomial for a particular choice of $v_{I}$, while Goldin–Singh [17] gave a type-independent proof to show that this formula works for all Coxeter elements.
@@ 3. Structure constants of Peterson Schubert calculus
For a general Dynkin diagram $I$, notice that $\sigma_{v_{I}}$, $\operatorname{det}\left(C_{I}\right)$, $W_{I}$, and $R\left(v_{I}\right)$ all factor multiplicatively with respect to the decomposition of connected components of $I$, hence the conclusion follows from the connected case. The proof is completed. ∎
@@
By letting $t=0$, the above arguments also give a direct proof of [21, Theorem 1.1.(3)] without going through the mixed $\Phi$-Eulerian numbers.
@@
on the Peterson variety $\operatorname{Pet}_{G}$. This vector bundle admits a well-defined section; see [3, Section 3.2] and [4, Section 5.2]. The zero locus of this section is exactly $\Omega_{I}:=\operatorname{Pet}_{G}\cap\Omega_{w_{I}}$ defined in (2.5); see [3, Proposition 3.12] and [4, Section 5.4]. Since $\operatorname{Pet}_{G,I}\cap\Omega_{I}=\left\{w_{I}\right\}$, it is not hard to see that $\left\{\varpi_{I}\right\}_{I\subset\Delta}$ is dual to the basis $\left\{\left[\operatorname{Pet}_{G,I}\right]\right\}_{I\subset\Delta}$ of $H^{S}_{*}(\operatorname{Pet}_{G};\mathbb{Z})$ up to a constant. This constant can be computed via integration over the Peterson Schubert variety $\operatorname{Pet}_{G,I}$, which is actually a product of smaller Peterson varieties; see [3, Section 4.2] for the computations in type $A$. This would give a different proof of the above theorem.
@@
$(B_{K}^{\widehat{s}})^{m}]_{i,k}=b_{i,k}^{J,m}=\sum_{a_{1}\in J\setminus\{i\}}\sum_{a_{2}\in J\setminus\{a_{1}\}}\cdots\sum_{a_{m-1}\in J\setminus\{a_{m-2}\}}b_{i,a_{1}}b_{a_{1},a_{2}}\cdots b_{a_{m-1},k}$
@@
It is not hard to see that we suffice to prove the lemma for connected $K$, since otherwise the matrix $B_{K}^{\widehat{s}}$ is block diagonal. Now suppose $K$ is a connected subset of $\Delta$, it is shown in [11, Lemma 2] that the eigenvalues of the Coxeter adjacency matrix $A_{K}$ are
@@
Similar to the proof of Lemma 3.7, the series $\sum\limits_{n=0}^{\infty}(B_{J})^{n}$ converges and thus
@@
where the Kronecker delta $\delta_{is}=0$ since $i\neq s$. The proof is completed. ∎
@@
Lusztig–Tits [30] gave an explicit formula for the $(i,j)$-entry of a Cartan matrix $C_{K}^{-1}$. If $i$ and $j$ are in different connected components of $K$, then $C_{K}^{-1}]_{i,j}=0$. Now assume $i$ and $j$ are in the same connected component of $K$. Let $i_{1},i_{2},\ldots,i_{p}$ be the unique sequence in $K$ such that $i_{1}=i,i_{p}=j$ and any two consecutive terms of the corresponding simple reflections $s_{i_{1}},s_{i_{2}},\ldots s_{i_{p}}$ are non-commuting. Let $W(i,j)$ be the parabolic subgroup of $W_{K}$ generated by $\left\{\sigma_{k};k\in K-\left\{i_{1},i_{2},\ldots,i_{p}\right\}\right\}$ and let $C(i,j)$ be the corresponding Cartan matrix. Then we have $C_{K}^{-1}]_{i,j}=\frac{\det(C(i,j))}{det(C_{K})}$.
@@
From the matrix we can see $d_{\{2\},\{1,2\}}^{\{1,2\}}=2t$, $d_{\{2\},\{1,2\}}^{\{1,2,3\}}=\frac{4}{3}$, which means $\varpi_{2}\cdot\varpi_{1}\varpi_{2}=2t\varpi_{1}\varpi_{2}+\frac{4}{3}\varpi_{1}\varpi_{2}\varpi_{3}.$ This matches the computation in [17, Example 6.8]
@@
$=$ $3456p_{\{1,2,3,4,5,6,7,8\}}+24p_{\{1,2,3,5,6,7,8,9\}}+240p_{\{1,3,4,5,6,7,8,9\}}$
which matches the computation for the non-equivariant case in [3, Example 5.2].
We have the following corollary. In particular, it means that our formulas provide an algebraic proof of the positivity of the equivariant structure constants of Peterson Schubert calculus.
@@
We have $d_{I,J}^{K}\geq 0$ for all $I,J,K\subset\Delta$.
@@
It is well known that the entries of the inverse of any Cartan matrix are all non-negative (see, for example, [22, Exercise 13.8]), hence by (3.4) and Lemma 3.8, all matrices $(d^{K}_{i,J})_{J,K}$ in Theorem 3.5 are polynomials in $t$ with non-negative coefficients. Since the entries in the product of such matrices remain polynomials in $t$ with non-negative coefficients, this corollary follows directly. ∎
@@
It follows that $K\supset I\cup J$. Let $K=K_{1}\bigsqcup\dots\bigsqcup K_{t}$ be the decomposition of $K$ into connected components. Without loss of generality, we assume $i\in K_{1}$ and $K\setminus K^{\prime}=\{s\}$ (here $s$ may or may not equal $i$). Note that $K^{\prime}\supset I^{\prime}\cup J$, we have $s\notin I^{\prime}$ and $s\notin J$. Since $d_{i,K^{\prime}}^{K}\neq 0$, by the proof of the $|I|=1$ case, it is not hard to see $s\in K_{1}$.
Assume $K_{1}\setminus\{s\}=K_{1,1}\bigsqcup\dots\bigsqcup K_{1,r}$ is the decomposition of $K_{1}\setminus\{s\}$ into connected components. Here $r=0$ means $K_{1}=\{i\}=\{s\}$, and $r=1$ means $K_{1}\setminus\{s\}$ is still connected.
@@
As for the sufficiency, assume $K\supset I\cup J$, and for each connected component $K_{k}$ of $K$, we have
@@
Applying the induction hypothesis to $I^{\prime}$, we have $d_{I^{\prime},J}^{K^{\prime}}\neq 0.$ This finishes the proof of the sufficiency. The proof is completed. ∎
@@ 4. Applications to mixed Eulerian numbers
We would like to introduce the mixed $\Phi$-Eulerian numbers for arbitrary Lie types as in [33] and their connections to the structure constants of the Peterson Schubert calculus. We mainly follow the notations as in [21].
@@
It is at most (n1)(n-1)-dimensional, sitting inside an affine hyperplane in $\mathbb{R}^{n}$. If $a_{1},a_{2},\ldots,a_{n}$ are distinct, then $P_{n}\left(a_{1},\ldots,a_{n}\right)$ is (n1)(n-1)-dimensional, and Postnikov gave a formula [33, Theorem 3.1] of its volume using divided symmetrization. This formula can be interpreted as a computation in the equivariant cohomology of the type $A$ permutohedral variety using the localization technique; see [21, Section 3].
@@
By this definition, the mixed $\Phi$-Eulerian number $A_{c_{1},\cdots,c_{n}}^{\Phi}$ is exactly the mixed volume of $c_{1}$ copies of $P_{\Phi}\left(\varpi_{1}\right)$, $c_{2}$ copies of $P_{\Phi}\left(\varpi_{2}\right),\cdots$, and $c_{n}$ copies of $P_{\Phi}\left(\varpi_{n}\right)$, multiplied by $n!$. Here, the weight polytopes $P_{\Phi}\left(\varpi_{1}\right),P_{\Phi}\left(\varpi_{2}\right),\ldots,P_{\Phi}\left(\varpi_{n}\right)$ are called the $\Phi$-hypersimplices. The mixed $\Phi$-Eulerian numbers are known to be non-negative integers; see [33] for more details. When $\Phi$ is of type $A$, these numbers are simply called mixed Eulerian numbers.
Postnikov provided in [33] a combinatorial formula for the mixed $\Phi$-Eulerian numbers in terms of certain binary trees. In [7], Berget–Spink–Tseng studied the log-concavity of matroid $h$-vectors in relation to mixed Eulerian numbers, using the fact that the cohomology ring of the type $A$ Permutohedral variety is exactly the Chow ring of the Boolean matroid. In [31], Nadeau–Tawari found a beautiful relation between mixed $\Phi$-Eulerian numbers and intersection numbers of Schubert varieties and the permutohedral variety for arbitrary Lie types. In [21], Horiguchi showed that mixed Eulerian numbers can be written as intersection numbers of Schubert divisors in Peterson variety for an arbitrary Lie type as follows.
Theorem 4.2 ([21, Theorem 1.1]).
@@
which is the volume polynomials of the nef divisors $\varpi_{1},\cdots,\varpi_{n}$ on the Peterson variety $\operatorname{Pet}_{G}$. While the Peterson variety is singular in general [26, Theorem 6], its rational cohomology ring $H^{*}(\operatorname{Pet}_{G};\mathbb{Q})$ in (2.6) admits the structure of the cohomology ring of a (rational smooth) toric orbifold (see, for example, [4] and [18]), thus satisfying all the Kähler package—the Poincaré duality, the hard Lefschetz theorem, and the Hodge–Riemann bilinear relation. This gives a different proof that $\operatorname{Vol}P_{\Phi}\left(u_{1}\varpi_{1}+\cdots+u_{n}\varpi_{n}\right)$ is Lorentzian in the sense of [10].
Using the above theorem, Horiguchi gave a combinatorial model introduced in [3] for the computation of mixed Eulerian numbers, and derived a type-by-type computation for mixed $\Phi$-Eulerian numbers for general Lie types by iteratively applying the Monk formula of Drellich in [13]. As an application of our main theorem, we now derive a type-uniform formula for the mixed $\Phi$-Eulerian numbers in arbitrary Lie types.
@@
Let $\Phi$ be an irreducible root system. Let $c_{1},\ldots,c_{n}$ be non-negative integers with $c_{1}+\cdots+c_{n}=n$. The mixed $\Phi$}-Euler number $A_{c_{1}\cdots c_{n}}^{\Phi}$ can be computed using the following formula:
@@
The proof is completed. ∎
The table of explicit values of $\frac{|W_{\Phi}|}{\det(C_{\Phi})}$ is presented below. Now we give some examples using our formula to compute mixed $\Phi$-Eulerian numbers.
@@
which matches the computation in [21, Example 5.4].
@@
which matches the computation in [21, Example 7.21].
It can be seen from the above examples that our formula avoids the need to discuss the changes in the Lie types case-by-case when considering sub-root systems as in [21, Section 7]. The only thing needed is to substitute $(c_{1},\ldots,c_{n})$ into our formula to obtain the results.

### v2

@@ 2.3. Cohomology ring of Peterson varieties and Peterson Schubert classes
Consider the restriction of the line bundle $L_{\mu}$ on the Peterson variety $\operatorname{Pet}_{G}$, which will also be denoted by $L_{\mu}$ when there is no confusion. Harada–Horiguchi–Masuda [20] gave the following explicit and beautiful Borel-type presentation of the equivariant cohomology ring of the Peterson varieties in all Lie types.
@@
Suppose that rank of $G$ is $n$. The $S$-equivariant cohomology ring of the Peterson variety $\operatorname{Pet}_{G}$ has the following presentation
@@
It follows that while the class $\iota^{*}\left(\sigma_{v_{I}}\right)$ depends on the choice of $v_{I}$, the class $p_{I}$ is independent of this choice. In [16, Remark 7.7], a type-independent formula for the intersection multiplicity $m\left(v_{I}\right)$ was conjectured, which was proved in [17].
@@
Next, we recall the equivariant Giambelli formula for the Peterson variety in [13] and [17]. It is well known that the equivariant Schubert classes $\sigma_{s_{i}},i\in\Delta$, generate $H_{T}^{*}(G/B;\mathbb{Q})$ as a $H_{T}^{*}(pt)$-algebra. Since the restriction map $H_{T}^{*}\left(G/B;\mathbb{Q}\right)\rightarrow H_{S}^{*}\left(\operatorname{Pet}_{G};\mathbb{Q}\right)$ is surjective [20], it is easy to see that each $\iota^{*}\sigma_{v_{I}}$ can be expressed as a polynomial with $\mathbb{Q}[t]$-coefficients in $\iota^{*}\sigma_{s_{i}},i\in\Delta$. Drellich [13] gave a very simple equivariant Giambelli formula for this polynomial for a particular choice of $v_{I}$, while Goldin–Singh [17] gave a type-independent proof to show that this formula works for all Coxeter elements.
@@ 3. Structure constants of Peterson Schubert calculus
For a general Dynkin diagram $I$, notice that $\sigma_{v_{I}}$, $\operatorname{det}\left(C_{I}\right)$, $W_{I}$, and $R\left(v_{I}\right)$ all factor multiplicatively with respect to the decomposition of connected components of $I$, hence the conclusion follows from the connected case. ∎
@@
By letting $t=0$, the above arguments also give a direct proof of [22, Theorem 1.1.(3)] without going through the mixed $\Phi$-Eulerian numbers.
@@
on the Peterson variety $\operatorname{Pet}_{G}$. This vector bundle admits a well-defined section; see [2, Section 3.2] and [4, Section 5.2]. The zero locus of this section is exactly $\Omega_{I}:=\operatorname{Pet}_{G}\cap\Omega_{w_{I}}$ defined in (2.5); see [2, Proposition 3.12] and [4, Section 5.4]. Since $\operatorname{Pet}_{G,I}\cap\Omega_{I}=\left\{w_{I}\right\}$, it is not hard to see that $\left\{\varpi_{I}\right\}_{I\subset\Delta}$ is dual to the basis $\left\{\left[\operatorname{Pet}_{G,I}\right]\right\}_{I\subset\Delta}$ of $H^{S}_{*}(\operatorname{Pet}_{G};\mathbb{Z})$ up to a constant. This constant can be computed via integration over the Peterson Schubert variety $\operatorname{Pet}_{G,I}$, which is actually a product of smaller Peterson varieties; see [2, Section 4.2] for the computations in type $A$. This would give a different proof of the above theorem.
@@
$(B_{K}^{\widehat{s}})^{m}]_{i,k}=\sum_{a_{1}\in J\setminus\{i\}}\sum_{a_{2}\in J\setminus\{a_{1}\}}\cdots\sum_{a_{m-1}\in J\setminus\{a_{m-2}\}}b_{i,a_{1}}b_{a_{1},a_{2}}\cdots b_{a_{m-1},k}.$
@@
It suffices to prove the lemma for connected $K$, since otherwise the matrix $B_{K}^{\widehat{s}}$ is block diagonal. Now suppose $K$ is a connected subset of $\Delta$, it is shown in [11, Lemma 2] that the eigenvalues of the Coxeter adjacency matrix $A_{K}$ are
@@
Similar to the proof of Lemma 3.7, the series $\sum\limits_{n=0}^{\infty}(B_{J})^{n}$ converges and
@@
where the Kronecker delta $\delta_{is}=0$ since $i\neq s$. ∎
@@
Lusztig–Tits [31] gave an explicit formula for the $(i,j)$-entry of a Cartan matrix $C_{K}^{-1}$. If $i$ and $j$ are in different connected components of $K$, then $C_{K}^{-1}]_{i,j}=0$. Now assume that $i$ and $j$ are in the same connected component of $K$. Let $i_{1},i_{2},\ldots,i_{p}$ be the unique sequence in $K$ such that $i_{1}=i,i_{p}=j$ and any two consecutive terms of the corresponding simple reflections $s_{i_{1}},s_{i_{2}},\ldots s_{i_{p}}$ are non-commuting. Let $W(i,j)$ be the parabolic subgroup of $W_{K}$ generated by $\left\{\sigma_{k};k\in K-\left\{i_{1},i_{2},\ldots,i_{p}\right\}\right\}$ and let $C(i,j)$ be the corresponding Cartan matrix. Then we have $C_{K}^{-1}]_{i,j}=\frac{\det(C(i,j))}{det(C_{K})}$.
@@
From the matrix we can see $d_{\{2\},\{1,2\}}^{\{1,2\}}=2t$, $d_{\{2\},\{1,2\}}^{\{1,2,3\}}=\frac{4}{3}$, which means $\varpi_{2}\cdot\varpi_{\{1,2\}}=2t\varpi_{\{1,2\}}+\frac{4}{3}\varpi_{\{1,2,3\}}.$ This matches the computation in [17, Example 6.8]
@@
$=$ $3456p_{\{1,2,3,4,5,6,7,8\}}+24p_{\{1,2,3,5,6,7,8,9\}}+240p_{\{1,3,4,5,6,7,8,9\}},$
which matches the computation for the non-equivariant case in [2, Example 5.2].
We have the following corollary. In particular, it means that our formulas provide an algebraic proof of the positivity of the equivariant structure constants of the Peterson Schubert calculus.
@@
The structure constant $d_{I,J}^{K}$ (and the structure constant $c_{I,J}^{K}$) is a polynomial in $t$ with non-negative coefficients for all $I,J,K\subset\Delta$.
@@
It is well known that the entries of the inverse of any Cartan matrix are all non-negative (see, for example, [23, Exercise 13.8]), hence by (3.4) and Lemma 3.8, all matrices $(d^{K}_{i,J})_{J,K}$ in Theorem 3.5 are polynomials in $t$ with non-negative coefficients. Since the entries of the resulting matrix product are still polynomials in $t$ with non-negative coefficients, this corollary follows directly. ∎
@@
It follows that $K\supset I\cup J$. Let $K=K_{1}\bigsqcup\dots\bigsqcup K_{t}$ be the decomposition of $K$ into connected components. Without loss of generality, we assume that $i\in K_{1}$ and $K\setminus K^{\prime}=\{s\}$ (here $s$ may or may not equal $i$). Note that $K^{\prime}\supset I^{\prime}\cup J$, we have $s\notin I^{\prime}$ and $s\notin J$. Since $d_{i,K^{\prime}}^{K}\neq 0$, by the proof of the $|I|=1$ case, it is not hard to see $s\in K_{1}$.
Assume that $K_{1}\setminus\{s\}=K_{1,1}\bigsqcup\dots\bigsqcup K_{1,r}$ is the decomposition of $K_{1}\setminus\{s\}$ into connected components. Here $r=0$ means $K_{1}=\{i\}=\{s\}$, and $r=1$ means $K_{1}\setminus\{s\}$ is still connected.
@@
As for the sufficiency, assume that $K\supset I\cup J$, and for each connected component $K_{k}$ of $K$, we have
@@
Applying the induction hypothesis to $I^{\prime}$, we have $d_{I^{\prime},J}^{K^{\prime}}\neq 0.$ This finishes the proof of the sufficiency. ∎
@@ 4. Applications to mixed Eulerian numbers
Firstly, we introduce the mixed $\Phi$-Eulerian numbers for arbitrary Lie types as in [34] and their connections to the structure constants of the Peterson Schubert calculus. We mainly follow the notation as in [22].
@@
It is at most (n1)(n-1)-dimensional, sitting inside an affine hyperplane in $\mathbb{R}^{n}$. If $a_{1},a_{2},\ldots,a_{n}$ are distinct, then $P_{n}\left(a_{1},\ldots,a_{n}\right)$ is (n1)(n-1)-dimensional, and Postnikov gave a formula [34, Theorem 3.1] of its volume using divided symmetrization. This formula can be interpreted as a computation in the equivariant cohomology of the type $A$ permutohedral variety using the localization technique; see [22, Section 3].
@@
By this definition, the mixed $\Phi$-Eulerian number $A_{c_{1},\ldots,c_{n}}^{\Phi}$ is exactly the mixed volume of $c_{1}$ copies of $P_{\Phi}\left(\varpi_{1}\right)$, $c_{2}$ copies of $P_{\Phi}\left(\varpi_{2}\right),\ldots$, and $c_{n}$ copies of $P_{\Phi}\left(\varpi_{n}\right)$, multiplied by $n!$. Here, the weight polytopes $P_{\Phi}\left(\varpi_{1}\right),P_{\Phi}\left(\varpi_{2}\right),\ldots,P_{\Phi}\left(\varpi_{n}\right)$ are called the $\Phi$-hypersimplices. The mixed $\Phi$-Eulerian numbers are known to be non-negative integers; see [34] for more details. When $\Phi$ is of type $A$, these numbers are simply called the mixed Eulerian numbers.
Postnikov provided in [34] a combinatorial formula for the mixed $\Phi$-Eulerian numbers in terms of certain binary trees. In [7], Berget–Spink–Tseng studied the log-concavity of matroid $h$-vectors in relation to the mixed Eulerian numbers, using the fact that the cohomology ring of the type $A$ Permutohedral variety is exactly the Chow ring of the Boolean matroid. In [32], Nadeau–Tawari found a beautiful relation between the mixed $\Phi$-Eulerian numbers and intersection numbers of Schubert varieties and the permutohedral variety for arbitrary Lie types. In [22], Horiguchi showed that the mixed Eulerian numbers can be written as intersection numbers of Schubert divisors in the Peterson variety for an arbitrary Lie type as follows.
Theorem 4.2 ([22, Theorem 1.1]).
@@
which is the volume polynomials of the nef divisors $\varpi_{1},\ldots,\varpi_{n}$ on the Peterson variety $\operatorname{Pet}_{G}$. While the Peterson variety is singular in general [27, Theorem 6], its rational cohomology ring $H^{*}(\operatorname{Pet}_{G};\mathbb{Q})$ in (2.6) admits the structure of the cohomology ring of a (rational smooth) toric orbifold (see, for example, [4] and [19]), which satisfies all the Kähler package—the Poincaré duality, the hard Lefschetz theorem, and the Hodge–Riemann bilinear relation. This gives a different proof that $\operatorname{Vol}P_{\Phi}\left(u_{1}\varpi_{1}+\cdots+u_{n}\varpi_{n}\right)$ is Lorentzian in the sense of [10].
Using the above theorem, Horiguchi gave a combinatorial model introduced in [2] for the computation of the mixed Eulerian numbers, and derived a type-by-type computation for the mixed $\Phi$-Eulerian numbers for general Lie types by iteratively applying the Monk formula of Drellich in [13]. As an application of our main theorem, we now derive a type-uniform formula for the mixed $\Phi$-Eulerian numbers in arbitrary Lie types.
@@
Let $\Phi$ be an irreducible root system. Let $c_{1},\ldots,c_{n}$ be non-negative integers with $c_{1}+\cdots+c_{n}=n$. The mixed $\Phi$}-Eulerian number $A_{c_{1},\ldots,c_{n}}^{\Phi}$ can be computed using the following formula:
@@
The table of explicit values of $\frac{|W_{\Phi}|}{\det(C_{\Phi})}$ is presented below. Now we give some examples using our formula to compute the mixed $\Phi$-Eulerian numbers.
@@
which matches the computation in [22, Example 5.4].
@@
which matches the computation in [22, Example 7.21].
It can be seen from the above examples that our formula avoids the need to discuss the changes in the Lie types case-by-case when considering sub-root systems as in [22, Section 7]. The only thing needed is to substitute $(c_{1},\ldots,c_{n})$ into our formula to obtain the results.